The sum of three consecutive natural numbers is 15 more than the greatest of them. What is the greatest of the three numbers?
Answer by mananth(16949)

let the numbers be x,x+1,x+2
...
x+x+1+x+2=15+x+2
3x+3=x+17
-x
3x-x+3=x-x+17
2x+3=17
-3
2x-3+3 = 17-3
2x=14
/2
x= 7
the third number ix 7+2=9